What is the mass of a piece of aluminum that starts at 100oC, is immersed in a calorimeter with 200 g of water at 25.0oC, and causes the water to rise to a temperature of 29.5oC?

200.0 g

53.4 g

59.3 g

4,184.0 g
Answer #1

m(water) = 200.0 g

T(water) = 25.0 oC

C(water) = 4.184 J/goC

m(metal) = to be calculated

C(metal) = 0.902 J/goC

T = 29.5 oC

We will be using heat conservation equation

use:

heat lost by metal = heat gained by water

m(metal)*C(metal)*(T(metal)-T) = m(water)*C(water)*(T-T(water))

m(metal)*0.902*(100.0-29.5) = 200.0*4.184*(29.5-25.0)

m(metal)*63.591 = 3765.6

m(metal)= 59.3 g

Answer: 59.3 g